A price decreased from $36 to $27. What is the percent of decrease?
My answer was 3.24% but I dont think that this is the right answer.
Thanks

Answer by CharStar(110)

The formula for percent change is: percent change = change/starting price
$36-$27 = $9
9/36 = 1/4 = 2.5%
I hope that helps you.